# Forms of Understanding for XAI-Explanations

### v2

@@ 3.3 Deep comprehension, shallow enabledness
One reason for deep understanding and shallow enabling is what has been called ‘inert knowledge’ [93, p. 5], although it was more concerned with inert (scientific) ideas than inert knowledge of individuals. More recent psychological research has demonstrated the prevalence and relevance of this problem and suggested several explanations for it that are different in nature and underscore its complexity [79]. The relationship between deep understanding and shallow enabledness is here conceptualized in a very specific way. According to meta-process explanations, the knowledge would be applicable, i.e., would result in enabledness, but external prerequisites such as motivation or the perception of benefits of applying the knowledge are missing. Also, persons might lack information about when or where to apply knowledge, which would also show up as shallow enabledness. Other explanations focus on a deficiency in the structure of knowledge. Important ones that could also explain the combination of deep understanding and shallow explanations are a lack of compilation and a lack conditionalized knowledge. Declarative knowledge has to be compiled in order to be used, and if compilation is insufficient or if a person lacks knowledge about conditions in which knowledge can or cannot be applied, shallow enabledness results. Finally, it should be noted that theories of situated cognition do not regard inert knowledge as a problem, but rather as due to the fact that knowledge is bound to certain contexts and situations and their specific conditions [e.g., 44].
@@ 3.4 Deep comprehension, deep enabledness
Agency is defined in various ways, which is apparent as the term itself is often combined with another term to form an open compounds part of a specific phrase such as, for example, ‘digital agency’ (Arnold and Clarke 7, p. 736; see, e.g., Passey et al. 72). There is the need for scholars to provide a precise definition of the term [cf. 1, p. 130] and the vagueness of the term is criticized, too [34]. Different definitions help to illustrate the problem: Passey et al. [72] define ‘digital agency’ quite broadly as “Digital Agency (DA)—consisting of digital competence, digital confidence, and digital accountability—is the individual’s ability to control and adapt to a digital world”. [27], in his inaugural lecture, was more explicit: First, agency, by which I mean not brute acts (of clicking on this button, pressing ‘like’ to this post) but (following Weber) the longer processes of action based on reflection, giving an account of what one has done, even more basically, making sense of the world so as to act within it. Kallinikos [52, p. 289f] describes how technology may influence agency “by inviting specific courses of action” that may be shaped to a certain degree (‘malleability’) and, that agency is “captured by the distinctive ways by which a technology invites people to frame a delimited domain of tasks or activities and organize their execution”. Arnold and Clarke [7, p. 751] define student agency as “[t]he discursive practice of positioning oneself or being positioned as responsible”. Emirbayer and Mische [34, p. 963] follow the pragmatic tradition of Bourdieu and Giddens, who define “[…] human agency as habitual, repetitive, and taken for granted—a view shared by ethnomethodologists, new institutionalists in organizational theory, and many others”. According to Sharma [85, p. 300] agency is a “contingently emergent feature of situated local action”. Thus, humans do not just have agency, but they can show agency through their actions. Siry and Lang [87] describe agency as a feature a person can acquire and develop. This means humans may acquire agency during childhood and education. Basu [11, p. 891] describes agency as purposefully examining and reflecting the actions of one person; the foundation of this acting is the beliefs and goals one person has. She says: “[…] I describe agency as purposefully considering and enacting both small- and large-scale change in personal and community domains, based on one’s beliefs and goals”. Duranti [31, p. 453] defines agency more broadly, as something entities are having. His working definition defines agency as a property of entities to (i) have a certain degree of control over their behavior, (ii) whose actions have a certain effect on other entities, and (iii) whose actions can be evaluated for example for responsibility.
@@
A mid to high degree of enabledness combined with a mid to high degree of comprehension are needed for agency. If both are pronounced, an individual is capable of actively manipulating and controlling various situations flexibly, by using what they know to reach individual goals. A high degree of comprehension and enabling is agency as it is a prerequisite for various properties mentioned in definitions by Passey et al. [72] (“control and adapt”), Basu [11] (“purposefully considering and enacting both small- and large-scale change in personal and community domains, based on one’s beliefs and goals”), Duranti [31] (having a certain degree of control over one’s own behavior, having an effect on other entities through one’s behaviour) and Sharma [85] (“a contingently emergent feature of situated local action” perceivable within actions”).
The question of how one would assess understanding in specific contexts and therefore, implicitly, how agency arises, is challenging is the fact that our concept of understanding remains a latent construct that is not trivial to define. Especially mental operations like connecting new facts cannot easily be measured. In other words, understanding cannot be assessed by asking to recall facts. However, it can be observed when people are engaged in unknown situations that require the utilization of previously acquired knowledge and a person’s enablement. In these situations, agency results in a flexible solution to the given task. This approach adheres to Perkins’ idea of measuring the ‘performance of understanding’ [73, 74]. Perkins recognizes “understanding through a flexible performance criterion” that “shows its face when people can think and act flexibly around what they know”. In contrast to this, a signal of lack of understanding is, “when a learner cannot go beyond rote and routine thought and action” [74, p. 42].
@@ 4 Dynamics of Understanding
Understanding is a process that unfolds over time when an explanation is successful [53]. In this section we offer a conceptualization of the dynamics of understanding over time. We focus on two distinct aspects that were mentioned in passing above: differences between novices and experts, and the intertwining of comprehension and enabledness in the transition from a lower to a higher form of understanding.
@@ 4.1 Novices and experts
First, novices and experts are assumed to differ in terms of procedural and conceptual knowledge. Experts were found to categorize problems by applying abstract solution procedures (e.g., major physical principles), whereas novices build their problem categorization on “the entities contained in the problem statement” [21, p. 15]. In terms of problem representation, experts translate features of the problem, thereby activating abstract category schemes for a problem type that are organized, for example, by universal laws or principles of nature. In contrast, novices organize their representation of a problem by schemata for object categories, thus staying on more concrete levels when formulating the problem. However, a novice can become an expert in a field by acquiring the relevant domain knowledge. The concept of the beginner is therefore fundamentally development-oriented.
@@
Second, while domain knowledge is one aspect in which novices and experts undoubtedly differ, they are also likely to differ in terms of experience, e.g., in a communicative practice. A study investigating novice tutors’ abilities to assess and monitor students’ knowledge was found to be profoundly inaccurate [23].
@@
In our attempt to go beyond classical cognitive psychology, which distinguishes between experts and novices by focusing on (domain) knowledge, we would like to enrich the pair of terms by bringing in constructive learning theory, which considers context orientation and situatedness as crucial. The focus here is on the interactive processes of how experts guide novices into ways of participating in a practice after a shared goal has been established in a given situation [61, 81]. For example, children are often referred to as novices because they are less knowledgeable. In conversations, the more knowledgeable participant (i.e., usually the adult) provides interactive support to help the child perform a aspired task [98]. Goodwin [41] broadens the perspective even further by explicitly including the body in the conceptual pair, thereby including multimodal aspects of interaction. He points out that in order to become competent, novices need to train both their bodies and the perceptually relevant structures of the objects or surroundings to be scrutinized in order to become skilled at performing a particular practice. In this sense, Goodwin [41] provides an anthropological view of embodied cognition that sees body and mind as inextricably intertwined. Understanding is therefore displayed through correct embodied behavior, and teachers could use displays of embodied behavior to assess students’ knowledge [see also 100, for perceptual structure provided by caregivers]. Hindmarsh et al. [49] shows how supervisors of dental students rely not so much on talk but on bodily behavior to assess their students’ ongoing understanding: even though students may claim to understand a procedure, the way they move their bodies to look into a patient’s mouth may actually reveal their lack of understanding of the matter at hand.
@@
Including embodiment in the conceptualization of understanding adds a new perspective because it may provide a more complete view of understanding than simply focusing on internal domains of knowledge. Novices would then differ from experts not only in terms of conceptualization and problem-solving procedures with respect to a particular domain, but also in that they would not have the embodied knowledge of how to perform a task [100]. It has been pointed out that knowledge and its asymmetries are relevant in distinguishing between experts and novices. The next section looks at different types of knowledge in this context.
@@ 4.2 Intertwinement of comprehension and enabledness: Transitions between the forms
The idea for the intertwinement of comprehension and enabledness (see Section 3.2) is related to the concept of ‘bootstrapping’ which might be an underlying mechanism facilitating the transgression from one part of the four-field diagram to another. Bootstrapping is assumed as being a learning mechanism for linguistic development by which growing abilities on one linguistic level, e.g., phonology, supports the progression on other linguistic levels such as syntax or semantics [68].
@@
This mechanism can be illustrated using the example 1 for the early stages of understanding (see Section 3.1). The challenge for a novice is to understand by explanation a stopping criterion of an iterative algorithm, or mainly to set up a strategy to understand the basic theory of a while-loop in a programming language. To describe the states of the model of a novice’s understanding in this example, we are looking at the transitions to other forms of understanding from the field of shallow comprehension and shallow enabledness:
@@ 5 Conclusion
As shown in the previous sections, discussing the cognitive side of XAI, it is important to consider types and forms of human understanding of explanations depending on their prior knowledge, motivation, and goals. These aspects are increasingly considered in the XAI literature.
We suggested aspects and forms of understanding which may occur in explanations between humans, but also in XAI-explanations for human users. Considering the forms given in our four-field diagram underlines that explainers as well as explainees pursue many different goals: to be enabled to do something, to comprehend something, or to achieve full agency. This agency needs – as we assume – both, deep comprehension and deep enabledness. There may, however, also be situations, in which humans only strive for shallow enabledness and/or shallow comprehension. The development of XAI technology has to consider these possibilities as well.
Generally, understanding is a journey from shallow to deep comprehension and enabledness, which results in both forms being highly intertwined and in need of being considered. In the end, it is not possible to reach deep enabledness without comprehension. For XAI, it follows that the amount of information required needs to be negotiated between the explainer (which could be an XAI system) and their explainee. As an important background of our contribution, Rohlfing et al. [82] conceptualize explaining as a social practice in which explainers and explainees are equally involved in (and responsible for) co-constructing the explanation and its understanding.
We consider these conceptualizations to be a concise theoretical structure for an interdisciplinary discussion and multi-perspectival development and evaluation of XAI technology, and explaining and understanding more generally. Two issues that we have not addressed above but that are highly relevant for explainability are (1) the need to consider that humans often have a subjective understanding that does not fully align with their actual/objective understanding and (2) that our cognitive approach of conceptualizing understanding neglects that explanations always take place in a social context. We close this paper by discussing these two aspects.
@@
In XAI, methods and interactive explanation systems that aim at co-constructing the human user’s understanding, relying on monitoring and adaptation processes for tailoring explanations to users’ needs [82], need to be aware of the difference between subjective and objective understanding. Depending on the mode of interaction, evidence of understanding that users provide can simply be subjective claims of understanding, (e.g., through the production of verbal acknowledgments signals or indications of satisfaction in surveys) or demonstrations of understanding that are more revealing of their actual understanding (e.g., being able to execute specific actions or asking follow-up questions that are contingent on understanding; cf. Koole 57). XAI systems need to be able to identify these different types of evidence, weigh them accordingly, and assess them critically when monitoring and adapting to their human interaction partners.

### v3

@@ 3.3 Deep comprehension, shallow enabledness
One reason for deep understanding and shallow enabling is what has been called ‘inert knowledge’ (Whitehead, 1929, p. 5), although it was more concerned with inert (scientific) ideas than inert knowledge of individuals. More recent psychological research has demonstrated the prevalence and relevance of this problem and suggested several explanations for it that are different in nature and underscore its complexity (Renkl et al., 1996). The relationship between deep understanding and shallow enabledness is here conceptualized in a very specific way. According to meta-process explanations, the knowledge would be applicable, i.e., would result in enabledness, but external prerequisites such as motivation or the perception of benefits of applying the knowledge are missing. Also, persons might lack information about when or where to apply knowledge, which would also show up as shallow enabledness. Other explanations focus on a deficiency in the structure of knowledge. Important ones that could also explain the combination of deep understanding and shallow explanations are a lack of compilation and a lack conditionalized knowledge. Declarative knowledge has to be compiled in order to be used, and if compilation is insufficient or if a person lacks knowledge about conditions in which knowledge can or cannot be applied, shallow enabledness results. Finally, it should be noted that theories of situated cognition do not regard inert knowledge as a problem, but rather as due to the fact that knowledge is bound to certain contexts and situations and their specific conditions (e.g., Greeno, 1998).
@@ 3.4 Deep comprehension, deep enabledness
Agency is defined in various ways, which is apparent as the term itself is often combined with another term to form an open compounds part of a specific phrase such as, for example, ‘digital agency’ (Arnold and Clarke 2013, p. 736; see, e.g., Passey et al. 2018). There is the need for scholars to provide a precise definition of the term (cf. Ahearn, 2001, p. 130) and the vagueness of the term is criticized, too (Emirbayer and Mische, 1998). Different definitions help to illustrate the problem: Passey et al. (2018) define ‘digital agency’ quite broadly as “Digital Agency (DA)—consisting of digital competence, digital confidence, and digital accountability—is the individual’s ability to control and adapt to a digital world”. (Couldry, 2014), in his inaugural lecture, was more explicit: First, agency, by which I mean not brute acts (of clicking on this button, pressing ‘like’ to this post) but (following Weber) the longer processes of action based on reflection, giving an account of what one has done, even more basically, making sense of the world so as to act within it. Kallinikos (2002, p. 289f) describes how technology may influence agency “by inviting specific courses of action” that may be shaped to a certain degree (‘malleability’) and, that agency is “captured by the distinctive ways by which a technology invites people to frame a delimited domain of tasks or activities and organize their execution”. Arnold and Clarke (2013, p. 751) define student agency as “[t]he discursive practice of positioning oneself or being positioned as responsible”. Emirbayer and Mische (1998, p. 963) follow the pragmatic tradition of Bourdieu and Giddens, who define “[…] human agency as habitual, repetitive, and taken for granted—a view shared by ethnomethodologists, new institutionalists in organizational theory, and many others”. According to Sharma (2008, p. 300) agency is a “contingently emergent feature of situated local action”. Thus, humans do not just have agency, but they can show agency through their actions. Siry and Lang (2010) describe agency as a feature a person can acquire and develop. This means humans may acquire agency during childhood and education. Basu (2008, p. 891) describes agency as purposefully examining and reflecting the actions of one person; the foundation of this acting is the beliefs and goals one person has. She says: “[…] I describe agency as purposefully considering and enacting both small- and large-scale change in personal and community domains, based on one’s beliefs and goals”. Duranti (2005, p. 453) defines agency more broadly, as something entities are having. His working definition defines agency as a property of entities to (i) have a certain degree of control over their behavior, (ii) whose actions have a certain effect on other entities, and (iii) whose actions can be evaluated for example for responsibility.
@@
A mid to high degree of enabledness combined with a mid to high degree of comprehension are needed for agency. If both are pronounced, an individual is capable of actively manipulating and controlling various situations flexibly, by using what they know to reach individual goals. A high degree of comprehension and enabling is agency as it is a prerequisite for various properties mentioned in definitions by Passey et al. (2018) (“control and adapt”), Basu (2008) (“purposefully considering and enacting both small- and large-scale change in personal and community domains, based on one’s beliefs and goals”), Duranti (2005) (having a certain degree of control over one’s own behavior, having an effect on other entities through one’s behaviour) and Sharma (2008) (“a contingently emergent feature of situated local action” perceivable within actions”).
The question of how one would assess understanding in specific contexts and therefore, implicitly, how agency arises, is challenging is the fact that our concept of understanding remains a latent construct that is not trivial to define. Especially mental operations like connecting new facts cannot easily be measured. In other words, understanding cannot be assessed by asking to recall facts. However, it can be observed when people are engaged in unknown situations that require the utilization of previously acquired knowledge and a person’s enablement. In these situations, agency results in a flexible solution to the given task. This approach adheres to Perkins’ idea of measuring the ‘performance of understanding’ (Perkins, 1993, 1998). Perkins recognizes “understanding through a flexible performance criterion” that “shows its face when people can think and act flexibly around what they know”. In contrast to this, a signal of lack of understanding is, “when a learner cannot go beyond rote and routine thought and action” (Perkins, 1998, p. 42).
@@ 4 Dynamics of Understanding
Understanding is a process that unfolds over time when an explanation is successful (Keil, 2006). In this section we offer a conceptualization of the dynamics of understanding over time. We focus on two distinct aspects that were mentioned in passing above: differences between novices and experts, and the intertwining of comprehension and enabledness in the transition from a lower to a higher form of understanding.
@@ 4.1 Novices and experts
First, novices and experts are assumed to differ in terms of procedural and conceptual knowledge. Experts were found to categorize problems by applying abstract solution procedures (e.g., major physical principles), whereas novices build their problem categorization on “the entities contained in the problem statement” (Chi et al., 1981, p. 15). In terms of problem representation, experts translate features of the problem, thereby activating abstract category schemes for a problem type that are organized, for example, by universal laws or principles of nature. In contrast, novices organize their representation of a problem by schemata for object categories, thus staying on more concrete levels when formulating the problem. However, a novice can become an expert in a field by acquiring the relevant domain knowledge. The concept of the beginner is therefore fundamentally development-oriented.
@@
Second, while domain knowledge is one aspect in which novices and experts undoubtedly differ, they are also likely to differ in terms of experience, e.g., in a communicative practice. A study investigating novice tutors’ abilities to assess and monitor students’ knowledge was found to be profoundly inaccurate (Chi et al., 2004).
@@
In our attempt to go beyond classical cognitive psychology, which distinguishes between experts and novices by focusing on (domain) knowledge, we would like to enrich the pair of terms by bringing in constructive learning theory, which considers context orientation and situatedness as crucial. The focus here is on the interactive processes of how experts guide novices into ways of participating in a practice after a shared goal has been established in a given situation (Lave and Wenger, 1991; Rogoff, 1990). For example, children are often referred to as novices because they are less knowledgeable. In conversations, the more knowledgeable participant (i.e., usually the adult) provides interactive support to help the child perform a aspired task (Wood et al., 1976). Goodwin (2013) broadens the perspective even further by explicitly including the body in the conceptual pair, thereby including multimodal aspects of interaction. He points out that in order to become competent, novices need to train both their bodies and the perceptually relevant structures of the objects or surroundings to be scrutinized in order to become skilled at performing a particular practice. In this sense, Goodwin (2013) provides an anthropological view of embodied cognition that sees body and mind as inextricably intertwined. Understanding is therefore displayed through correct embodied behavior, and teachers could use displays of embodied behavior to assess students’ knowledge (see also Zukow-Goldring, 1996, for perceptual structure provided by caregivers). Hindmarsh et al. (2011) shows how supervisors of dental students rely not so much on talk but on bodily behavior to assess their students’ ongoing understanding: even though students may claim to understand a procedure, the way they move their bodies to look into a patient’s mouth may actually reveal their lack of understanding of the matter at hand.
@@
Including embodiment in the conceptualization of understanding adds a new perspective because it may provide a more complete view of understanding than simply focusing on internal domains of knowledge. Novices would then differ from experts not only in terms of conceptualization and problem-solving procedures with respect to a particular domain, but also in that they would not have the embodied knowledge of how to perform a task (Zukow-Goldring, 1996). It has been pointed out that knowledge and its asymmetries are relevant in distinguishing between experts and novices. The next section looks at different types of knowledge in this context.
@@ 4.2 Intertwinement of comprehension and enabledness: Transitions between the forms
The idea for the intertwinement of comprehension and enabledness (see Section 3.2) is related to the concept of ‘bootstrapping’ which might be an underlying mechanism facilitating the transgression from one part of the four-field diagram to another. Bootstrapping is assumed as being a learning mechanism for linguistic development by which growing abilities on one linguistic level, e.g., phonology, supports the progression on other linguistic levels such as syntax or semantics (Morgan and Demuth, 1996).
@@
Example 1 can be used to illustrate this mechanism in the early stages of understanding (see Section 3.1). The challenge for a novice is to understand the stopping criterion of an iterative algorithm through explanation, or to devise a strategy for understanding the basic theory of a while-loop in a programming language. To describe the states of the model of a novice’s understanding in this example, we are looking at the transitions to other forms of understanding from the field of shallow comprehension and shallow enabledness:
@@
5 Forms of Understanding in a Classic XAI Scenario
In Section 1, we argued that our four-field model of forms of understanding is applicable to explanations in both human interaction and human-explaining AI interaction. The explaining AI, however, was assumed to be a new form of a ‘social‘ explainable AI which is co-constructive and interactive (Rohlfing et al., 2021, 2025). This raises the question of whether the model would, in principle, also be useful in ‘classic’ XAI scenarios, e.g., where explanations take the form of counterfactuals. A problem with such XAI scenarios is that users who are interested in explanations of outcomes produced by AI systems already require expertise in interpreting the form of explanation that is provided by a system. For example, they need to be able to work with LIME explanations (Ribeiro et al., 2016), or know how to formulate contrastive questions and interpret counterfactual explanations (Miller, 2021, 2023) – and how these forms of explanation can be used to improve understanding of outcomes.
In this section we show that our four-field model of understanding can also be usefully applied to such scenarios by applying the model to the classic XAI scenario in which an automated system decides whether to grant a client a bank loan and under what conditions. Let’s assume that the client uses a bank’s website that offers an interface for customers interested in loans and that this interface offers counterfactual explanations. We assume that this client has relevant XAI experience, i.e., they know how to interpret counterfactual explanations and ask contrastive questions, but have no domain expertise. In other words, they do not know much about how bank loans are calculated and decided upon.
When the client starts to interact with the system, they can thus be assumed to have shallow comprehension and shallow enabledness on how the bank will make the decision. First, the system requests relevant information from the client, such as their demographics, income and credit score. It also asks for information on the loan they are seeking, e.g., the loan amount, term and interest rate, and how it will be used. By providing all this information, the client may already gain some comprehension that the decision and conditions are probably based on the specific parameters they provided. The system initially rejects the client’s request, prompting them to engage with the XAI component and ask contrastive questions about their income, the loan amount, loan term, and its intended use. This deepens the client’s comprehension of the decision procedure, e.g., that the input variables are not necessarily independent, and that making smaller changes to multiple variables may be more effective than making larger changes to a single variable in order to achieve a different outcome.
By experimenting with the input variables and posing further contrastive questions, the client can explore the system’s decision-making process in more depth, deepening their enabledness to predict the decision based on changes to their input. They can now better understand which types of loans they are eligible for and which personal factors they are in principle able to change to qualify for the loan they initially aimed for (see, e.g., Mindlin et al. 2024 for an experimental investigation in measuring users’ enabledness to predict income using a dialogue-based XAI system). This gives them agency and empowers them to make informed decisions about their financial situation.
It is important to note that, in classic XAI scenarios, the client’s ability to achieve understanding depends on two prerequisites: (1) knowledge of how to interpret the forms of explanation provided by the system (e.g., formulating contrastive questions and interpreting counterfactual explanations); and (2) knowledge of how to systematically make use of the system’s forms of explanation to actually facilitate understanding of the its outcomes. As a result, the responsibility for achieving understanding rests entirely with the client. An interactive and co-constructive social XAI system, in contrast, would adapt its explanation and scaffold the client’s understanding (Rohlfing et al., 2021), making it a more collaborative process.
@@ 6 Conclusion
As shown in the previous sections, discussing the cognitive side of XAI, it is important to consider types and forms of human understanding of explanations depending on their prior knowledge, motivation, and goals. These aspects are increasingly considered in the XAI literature (Miller, 2019, 2023; Speith et al., 2024; Rohlfing et al., 2021, 2025).
We suggested aspects and forms of understanding which may occur in explanations between humans, but also in XAI-explanations for human users. Considering the forms given in our four-field diagram (Figure 1) underlines that explainers as well as explainees pursue many different goals: to be enabled to do something, to comprehend something, or to achieve full agency. This agency needs – as we assume – both, deep comprehension and deep enabledness. There may, however, also be situations, in which humans only strive for shallow enabledness and/or shallow comprehension. The development of XAI technology has to consider these possibilities as well.
Generally, understanding is a journey from shallow to deep comprehension and enabledness, which results in both forms being highly intertwined and in need of being considered. In the end, it is not possible to reach deep enabledness without comprehension. For XAI, it follows that the amount of information required needs to be negotiated between the explainer (which could be an XAI system) and their explainee. As an important background of our contribution, Rohlfing et al. (2021) conceptualize explaining as a social practice in which explainers and explainees are equally involved in (and responsible for) co-constructing the explanation and its understanding.
We consider these conceptualizations to be a concise theoretical structure for an interdisciplinary discussion and multi-perspectival development and evaluation of (social) XAI technology, and explaining and understanding more generally. Two issues that we have not addressed above but that are highly relevant for explainability are (1) the need to consider that humans often have a subjective understanding that does not fully align with their actual/objective understanding and (2) that our cognitive approach of conceptualizing understanding neglects that explanations always take place in a social context. We close this paper by discussing these two aspects.
@@
In XAI, methods and interactive explanation systems that aim at co-constructing the human user’s understanding, relying on monitoring and adaptation processes for tailoring explanations to users’ needs (Rohlfing et al., 2021), need to be aware of the difference between subjective and objective understanding. Depending on the mode of interaction, evidence of understanding that users provide can simply be subjective claims of understanding, (e.g., through the production of verbal acknowledgments signals or indications of satisfaction in surveys) or demonstrations of understanding that are more revealing of their actual understanding (e.g., being able to execute specific actions or asking follow-up questions that are contingent on understanding; cf. Koole 2010). XAI systems need to be able to identify these different types of evidence, weigh them accordingly, and assess them critically when monitoring and adapting to their human interaction partners.